### v2

# CoLA-Flow Policy: Temporally Coherent Imitation Learning via Continuous Latent Action Flow Matching for Robotic Manipulation

Abstract
Learning long-horizon robotic manipulation requires jointly achieving expressive behavior modelling, real-time inference, and stable execution, which remains challenging for existing generative policies. Diffusion-based approaches offer strong modelling capacity but incur high inference latency, while flow matching enables fast one-step generation yet often suffers from unstable execution when operating directly in the raw action space. We propose CoLA-Flow Policy, a trajectory-level imitation learning framework that performs flow matching in a continuous latent action space. By encoding action sequences into temporally coherent latent trajectories and learning an explicit latent-space flow, CoLA-Flow Policy decouples global motion structure from low-level control noise, enabling smooth and reliable long-horizon execution. The framework further integrates geometry-aware point cloud conditioning and execution-time multimodal modulation, with visual cues as a representative modality evaluated in real-world settings for robust deployment. Experiments in simulation and on real robots show that CoLA-Flow Policy achieves near single-step inference, improves trajectory smoothness by up to 93.7% and task success by up to 25% over raw action-space flow baselines, while remaining significantly faster than diffusion-based policies.
I INTRODUCTION
Diffusion-based policies (DP) [5] have demonstrated strong performance in robotic visuomotor imitation learning, particularly for complex manipulation tasks in both simulation and real-world environments. By modelling multimodal action distributions conditioned on rich sensory observations such as images [5, 16] or point clouds [33, 25, 26], these methods enable expressive behavior modelling and generalization across diverse task variations.
More recently, flow matching [15, 4, 9] has emerged as an alternative generative paradigm for robotic control [35, 6, 1]. By formulating action generation as the solution of an ordinary differential equation (ODE), flow matching enables near one-step inference and substantially improves inference efficiency compared to diffusion-based approaches, making it a promising framework for real-time imitation learning.
Despite these advances, deploying generative policies on physical robots remains challenging. Diffusion policies incur high inference latency due to iterative denoising, while flow matching methods generate actions by integrating a predicted flow field, a process that amplifies modelling noise over long horizons and leads to jittery, unstable trajectories. As a result, existing approaches struggle to simultaneously achieve expressive behavior modelling, fast inference, and stable execution.
Several recent works have explored different strategies to alleviate these issues. DP3 [33] and iDP3 [32] improve perception efficiency through lightweight point cloud encoders, RDP [28] introduces a hierarchical Fast–Slow control structure for long-horizon behaviours, and Flow Policy [35] incorporates consistency constraints to enhance execution performance. However, our experimental results indicate that the fundamental trade-off between inference efficiency and execution stability has not yet been fully resolved.
To address this challenge, we propose CoLA-Flow Policy, a trajectory-level action generation framework that performs flow matching in a continuous latent action space under geometry-aware 3D scene conditioning. Instead of generating controls directly in the physical action space, our approach models and generates entire action trajectories in a structured latent space, enabling efficient one-step inference while maintaining temporal coherence for long-horizon manipulation.
The key advantage of CoLA-Flow Policy lies in decoupling global motion structure from low-level control noise. By operating on motion-level latent representations, the learned flow captures smooth and consistent trajectory evolution, reducing the amplification of modelling errors during generation and leading to more stable execution. In addition, the latent trajectories can be decoded under environment-dependent modulation, allowing the policy to incorporate multi-modal sensory cues—such as visual feedback from a wrist-mounted camera at execution time—without altering the underlying generative process. This design results in reliable long-horizon behaviors with improved smoothness and robustness.
Our main contributions are summarized as follows:
  • We introduce a trajectory-level latent flow matching framework that reconciles fast inference with smooth and stable robotic manipulation.
  • We show that performing flow matching in a continuous latent action space significantly reduces trajectory jitter and improves execution reliability.
  • We demonstrate through comprehensive simulation and real-world experiments that the proposed method achieves robust performance across diverse tasks, including improved tolerance to heterogeneous and suboptimal demonstrations.
Figure 1: Overall architecture of the proposed CoLA-Flow Policy. The system first encodes point cloud observations into geometry-aware scene features, then performs latent-space flow matching to generate temporally coherent latent action trajectories, which are finally decoded into executable control commands under visual conditioning.
II RELATED WORKS
II-A Diffusion and Flow-Based Policies for Robotics
Diffusion-based policies [13, 5, 33, 32, 26, 16] model visuomotor control as conditional trajectory generation and achieve strong performance in complex manipulation tasks. However, their reliance on iterative denoising leads to high inference latency and sensitivity to execution noise, limiting real-time deployment. Acceleration strategies such as DDIM [21] or hierarchical designs like RDP [28] alleviate but do not eliminate the need for multi-step inference, leaving the recursive nature of diffusion a fundamental bottleneck.
Flow-based generative models [15, 4] provide a more direct alternative by learning vector fields that define probability paths via ODEs. Flow Matching enables fast inference by directly regressing velocity fields, and recent consistency-based variants [35, 6, 1] further reduce sampling steps. Despite their efficiency, flow-based policies often exhibit trajectory oscillations when applied directly in raw action spaces, especially under high-dimensional actions and complex visual inputs.
Our work builds on consistency flow matching while addressing this limitation through trajectory-level latent modelling and geometry-aware conditioning.
II-B Visual Imitation Learning with 3D Representations
Early visual imitation learning methods rely primarily on 2D image observations [5, 8], processed by convolutional networks [12] or vision transformers [7]. Recent 3D-based approaches [33, 32, 25] incorporate depth or point clouds and improve geometric reasoning, but commonly face two challenges.
First, prediction-and-planning formulations [27, 34] estimate 3D keyframes or target poses, introducing additional optimization at test time and scaling poorly to high-dimensional control. Second, efficient point cloud encoding remains difficult: PointNet-style architectures [19] can be computationally expensive [25], while simplified encoders [33, 32] trade geometric sensitivity for speed.
In contrast, our work follows a policy-centric design that directly outputs continuous actions, integrating structured 3D geometric representations while maintaining real-time-capable inference.
II-C Trajectory-Level Modelling and Dexterous Manipulation
Dexterous manipulation has traditionally been addressed using reinforcement learning or task-specific pipelines [23, 34], often requiring extensive training or carefully engineered structures. Recent imitation learning methods model action sequences or chunks [5, 35], but may still suffer from temporal inconsistency under long-horizon, high-dimensional control.
Trajectory-level modelling improves long-horizon coherence and robustness to noisy or suboptimal demonstrations, with latent action representation playing a central role. Discrete latents via VQ-VAE [24] are widely adopted [2, 3] to enforce information bottlenecks, but may limit expressiveness. Recent studies show that appropriately structured continuous latent representations can achieve strong performance without sacrificing stability [10]. For example, CoMo [29] mitigates shortcut learning through temporal difference modelling, and LAOM [17] demonstrates the benefits of higher-capacity continuous bottlenecks.
These results motivate trajectory-level continuous latent action modelling as an effective mechanism for generating smooth and temporally coherent dexterous behaviours, which we adopt in this work.
III METHODS
We propose CoLA-Flow Policy, a modular imitation learning framework for long-horizon robotic manipulation. As illustrated in Fig. 1, the method first learns a temporally coherent latent action representation, which serves as a compact trajectory-level abstraction. A generative model then operates in this latent space to produce long-horizon latent trajectories under geometric scene conditioning, and the resulting latent actions are finally decoded into executable control commands with execution-time sensory modulation.
III-A Coherent Latent Action Representation
Long-horizon robotic manipulation requires action trajectories that are both expressive and execution-stable. Applying flow-based generative models directly in the raw action space often leads to numerical instability and high-frequency noise amplification during trajectory integration. We therefore introduce a trajectory-level latent action modeling framework that encodes action sequences into a continuous and temporally coherent latent trajectory space and decodes generated latent trajectories into executable controls, thereby reshaping the generation space prior to flow-based generative modeling. The overall architecture is illustrated in Fig. 2.
Temporally Continuous Latent Action Modeling
Instead of modeling individual control commands, we encode short-horizon action segments into latent variables that represent motion-level abstractions. Given an action trajectory $\mathbf{A}_{t:t+H-1}\in\mathbb{R}^{H\times d_{a}}$, we divide it into $K$ consecutive chunks $\{\mathbf{A}^{(1)},\ldots,\mathbf{A}^{(K)}\}$, each spanning $c$ time steps. Each chunk is mapped to a compact feature embedding and processed by a recurrent encoder with gated memory. Specifically, we employ a GRU-based encoder to aggregate action chunks over time,
$\mathbf{h}_{k}=\mathrm{GRU}(\mathbf{x}_{k},\mathbf{h}_{k-1}),$ (1)
where $\mathbf{x}_{k}$ denotes the feature embedding of the $k$-th action chunk, obtained via a lightweight temporal convolution.
By conditioning each latent code on the entire history of preceding action segments, the recurrent encoder induces a continuous and history-dependent latent trajectory, introducing a strong temporal inductive bias that suppresses high-frequency variations. As a result, the latent action space is shaped into a smooth and temporally continuous trajectory manifold rather than a collection of independent per-chunk representations, which is critical for stable ODE-based flow matching. In contrast, non-recurrent per-chunk encoders break this continuity and empirically lead to increased trajectory jitter.
Figure 2: Trajectory-level latent action representation with recurrent encoding and conditional decoding. A GRU-based encoder introduces a temporal inductive bias that promotes coherent latent trajectories, while an MLP decoder modulated by wrist-camera features via FiLM enables visually adaptive action execution.
Variational Regularization
To further regularize the latent space, we adopt a variational formulation [14]. A KL divergence term between the posterior and a standard normal prior encourages compactness and continuity of the latent distribution, improving robustness to noisy demonstrations. This regularization operates at the trajectory level, making it well suited for long-horizon manipulation.
Visual-Conditioned Action Decoding
The decoder is designed to incorporate execution-time sensory information without interfering with latent trajectory generation. In principle, diverse modalities(e.g., vision, tactile, proprioception) can be integrated through feature-wise linear modulation (FiLM) [18], enabling flexible multimodal conditioning at low computational cost.
In this work, we instantiate this design using wrist-mounted camera observations, which are encoded by a lightweight visual encoder based on a pretrained ResNet-18. [11]. Execution-time images are processed by a lightweight visual encoder and injected into the decoder via FiLM layers. Given a latent action code $\mathbf{z}_{k}$, the decoder reconstructs the corresponding action chunk as
$\hat{\mathbf{A}}^{(k)}=f_{\theta}(\mathbf{z}_{k}\mid\mathbf{v}_{t}),$ (2)
where $\mathbf{v}_{t}$ denotes the encoded visual feature. By confining perceptual conditioning to the decoding stage, the framework cleanly separates latent trajectory planning from execution-time adaptation while preserving temporal coherence.
Training Objective and Protocol
The latent action encoder–decoder is first trained using a variational objective with reconstruction, KL regularization, and a lightweight smoothness constraint to learn a temporally coherent latent space. The learned representation is then fixed and used as a plug-in interface, upon which a downstream generative model is trained to generate long-horizon latent trajectories.
III-B Latent-Space Action Generation via Flow Matching
Given a sequence of latent action codes $\{\mathbf{z}_{1},\ldots,\mathbf{z}_{K}\}$ obtained from the trajectory-level encoder, we generate long-horizon manipulation behaviors by modelling their evolution in continuous time using flow matching [15]. Unlike diffusion-based policies [5] that rely on iterative denoising, flow matching enables efficient one-step generation. However, when applied directly in the raw action space, flow-based models are highly sensitive to local inconsistencies and noise. We therefore perform flow matching exclusively in the temporally coherent latent action space, which is explicitly designed to be smooth and temporally coherent (Sec. III-A).
Latent-Space Flow Dynamics
Let $\mathbf{z}$ denote a latent action code sampled from the target latent distribution and $\tilde{\mathbf{z}}$ a corresponding sample from a simple base distribution. We learn a time-dependent vector field $\boldsymbol{\nu}_{\theta}(t,\mathbf{z})$ that defines a continuous probability path from the source to the target distribution, governed by the ordinary differential equation
$\frac{d\xi_{\mathbf{z}}(t)}{dt}=\boldsymbol{\nu}_{\theta}(t,\xi_{\mathbf{z}}(t)),\quad\xi_{\mathbf{z}}(0)=\tilde{\mathbf{z}}$. (3)
Operating on motion-level latent representations rather than raw control signals substantially improves numerical stability, as the smooth geometry of the latent space suppresses the amplification of high-frequency variations during trajectory generation.
Consistency Flow Matching in Latent Space
To efficiently learn the latent vector field across noise levels, we adopt consistency flow matching (CFM) [30] and define the flow function
$f_{\theta}(t,\mathbf{z})=\mathbf{z}+(1-t)\,\boldsymbol{\nu}_{\theta}(t,\mathbf{z}),$ (4)
where $t\in[0,1]$ interpolates between the source and target latent distributions. Here, $\boldsymbol{\nu}_{\theta}(t,\mathbf{z})$ represents a learned velocity field in the latent action space, and the factor (1t)(1-t) attenuates update magnitudes as the trajectory approaches the target distribution, improving stability near convergence. We further apply a time-dependent input normalization
$c_{\mathrm{in}}(t)=\frac{1}{\sqrt{t^{2}+(1-t)^{2}}},$ (5)
scaling the latent state before feeding it into the velocity field network: $\boldsymbol{\nu}_{\theta}\!\left(t,\,c_{\mathrm{in}}(t)\cdot\mathbf{z}\right)$. This normalization balances latent state scales across noise levels and prevents overfitting to specific time regions.
One-Step Latent Trajectory Generation
At inference time, latent action codes are sampled from the base distribution and transformed via a single evaluation of the learned flow:
$\hat{\mathbf{z}}_{k}=f_{\theta}(1,\tilde{\mathbf{z}}_{k}).$ (6)
As with flow-based generative models, latent trajectories are obtained through a single-step transformation defined by the learned latent-space flow field, without requiring iterative denoising as in diffusion-based policies [5]. This enables one-shot generation of the entire latent action trajectory, avoiding iterative sampling and substantially reducing inference latency.
The resulting latent trajectory preserves temporal coherence by construction and is decoded into executable action segments using the multimodal-conditioned decoder described in Sec. III-A. By decoupling trajectory generation from execution-level adaptation, the policy achieves both fast inference and stable, smooth manipulation behavior.
III-C Geometry-Aware 3D Scene Conditioning
Stable execution of latent action trajectories requires accurate awareness of scene geometry, as geometric errors can propagate across multiple decoded action segments over long horizons. We therefore condition the latent flow policy on a compact, geometry-aware 3D scene representation derived from point clouds, following prior point-cloud conditioned manipulation policies [33].
Point Cloud Encoding
Scene point clouds are reconstructed from depth images captured by a fixed global camera, followed by workspace cropping to remove irrelevant background. Farthest Point Sampling (FPS) [19] is applied to select a sparse set of center points that uniformly cover the workspace, serving as anchors for local neighborhood construction.
To capture geometric information at multiple spatial scales, we adopt a dual-branch encoder consisting of a local encoder $f_{l}$ and a center encoder $f_{c}$, whose overall architecture is illustrated in Fig. 3. The local encoder aggregates relative point offsets within each neighborhood using a residual convolutional backbone with combined max–mean pooling, focusing on contact-level geometry that is invariant to global translation. In parallel, the center encoder maps the sampled center coordinates to a compact global context vector via a lightweight MLP, capturing coarse scene layout with minimal computational overhead:
$\mathbf{f}_{l}=f_{l}(\text{pcd}_{\text{local}}),\quad\mathbf{f}_{c}=f_{c}(\text{pcd}_{\text{center}}).$ (7)
Figure 3: Geometry-aware point cloud encoder. Local neighborhoods around farthest-sampled centers capture translation-invariant local geometry via residual convolutions and max–mean pooling, while a lightweight center encoder provides compact global scene context for latent trajectory conditioning.
Hierarchical Geometric Conditioning
The extracted geometric features are injected into the latent flow network through a two-stage FiLM mechanism. Given an intermediate activation $\mathbf{h}$, local geometric modulation is applied first, followed by center-level conditioning:
$\mathbf{h}^{\prime}$ $=\boldsymbol{\gamma}_{l}\odot\mathbf{h}+\boldsymbol{\beta}_{l},\quad[\boldsymbol{\gamma}_{l},\boldsymbol{\beta}_{l}]=\mathrm{MLP}_{l}(\mathbf{f}_{l}),$ (8)
$\mathbf{h}^{\prime\prime}$ $=\boldsymbol{\gamma}_{c}\odot\mathbf{h}^{\prime}+\boldsymbol{\beta}_{c},\quad[\boldsymbol{\gamma}_{c},\boldsymbol{\beta}_{c}]=\mathrm{MLP}_{c}(\mathbf{f}_{c}).$ (9)
Applying local modulation prior to center-level conditioning allows the policy to first adapt to contact-level constraints and subsequently adjust according to global scene structure, which we find to be more effective than the reverse ordering.
Overall, this geometry-aware conditioning provides the latent action model with robust spatial context, enabling stable long-horizon manipulation without compromising inference efficiency.
IV EXPERIMENTS
IV-A Evaluation Metrics
We evaluate all methods using three complementary metrics that jointly characterize task performance, real-time efficiency, and execution quality: task success rate, response time, and trajectory smoothness.
Task Success Rate.
Task success rate is defined as the percentage of trials that satisfy task-specific success criteria within the episode horizon. Success is determined by predefined geometric constraints, including object pose errors below a threshold and stable grasps or placements maintained for a minimum duration.
Response Time.
Response time measures the end-to-end inference latency from receiving sensory observations to outputting control actions, reflecting the real-time computational efficiency and closed-loop responsiveness of each policy.
Trajectory Smoothness.
To quantify long-horizon execution stability, we define a trajectory smoothness metric that captures both time-domain and frequency-domain characteristics of executed joint trajectories. Specifically, smoothness is computed as a weighted combination of discrete-time jerk and high-frequency spectral energy.
Given a sequence of joint actions $\{\mathbf{a}_{t}\}_{t=1}^{T}$ with control timestep $\Delta t$, jerk is approximated using a third-order finite difference:
$\mathbf{j}_{t}=\frac{\mathbf{a}_{t+1}-3\mathbf{a}_{t}+3\mathbf{a}_{t-1}-\mathbf{a}_{t-2}}{(\Delta t)^{3}}.$ (10)
The time-domain smoothness term is defined as the mean squared jerk magnitude:
$\mathcal{S}_{\mathrm{jerk}}=\frac{1}{T-2}\sum\nolimits_{t=3}^{T}\|\mathbf{j}_{t}\|_{2}^{2}.$ (11)
To further capture high-frequency oscillations not fully reflected by jerk alone, we compute the frequency-domain energy of each joint trajectory using the discrete Fourier transform (DFT). Let $\hat{\mathbf{a}}_{f}$ denote the Fourier coefficients of $\{\mathbf{a}_{t}\}$. The high-frequency energy ratio is defined as
$\mathcal{S}_{\mathrm{freq}}=\frac{\sum_{f>f_{c}}\|\hat{\mathbf{a}}_{f}\|_{2}^{2}}{\sum_{f}\|\hat{\mathbf{a}}_{f}\|_{2}^{2}},$ (12)
where $f_{c}$ is a cutoff frequency separating low- and high-frequency components.
The final trajectory smoothness metric is given by
$\mathcal{S}_{\mathrm{smooth}}=\alpha\,\mathcal{S}_{\mathrm{jerk}}+\beta\,\mathcal{S}_{\mathrm{freq}},$ (13)
where lower values indicate smoother trajectories with reduced abrupt variations and high-frequency oscillations. Based on extensive empirical evaluation and normalization of the two terms, we set $\alpha=0.25$ and $\beta=0.75$ in all experiments.
TABLE I: Main results for simulation experiments. Success rates (%) are reported as mean $\pm$ standard deviation. Avg. Success rate averages results over Adroit and MetaWorld. $\Delta$ Success / $\Delta$ Time denotes relative changes w.r.t. DP3 in average success rate and per-step inference latency.
Methods Success rate in Adroit Success rate in MetaWorld Avg. Success rate $\boldsymbol{\Delta}$ Success Avg. Time $\boldsymbol{\Delta}$ Time
Diffusion Policy $26.0\pm 4.0$ $37.3\pm 2.3$ $31.7\pm 6.9$ 34.0-34.0 $35.7\text{ ms}$ 20.2-20.2
Flow Policy $56.7\pm 2.3$ $65.3\pm 2.3$ $61.0\pm 5.2$ 4.7-4.7 $\mathbf{6.1\text{ ms}}$ $\mathbf{-49.8}$
DP3 $62.0\pm 2.0$ $69.3\pm 3.1$ $65.7\pm 4.6$ $0.0$ $55.9\text{ ms}$ $0.0$
iDP3 $62.7\pm 2.3$ $71.3\pm 2.3$ $67.0\pm 5.2$ +1.3+1.3 $58.9\text{ ms}$ +3.0+3.0
RDP $70.0\pm 2.0$ $76.0\pm 2.0$ $0\pm 3.7 +7.3+7.3$ $57.3\text{ ms}$ +1.4+1.4
CoLA-Flow Policy (Ours) $\mathbf{76.0\pm 2.0}$ $\mathbf{80.7\pm 3.1}$ $\mathbf{78.3\pm 3.4}$ $\mathbf{+12.6}$ $7.5\text{ ms}$ 48.4-48.4
IV-B Simulation Experiments
IV-B1 Simulation Setup
Simulation Benchmarks.
We evaluate all methods on two widely used manipulation benchmarks, Adroit [20] and Meta-World [31], covering 37 tasks in total. Adroit focuses on contact-rich dexterous manipulation with a high-dimensional articulated hand, while Meta-World consists of diverse arm-based manipulation tasks with varying difficulty. Both benchmarks are built upon the MuJoCo [22] physics simulator. For both benchmarks, expert demonstrations are collected using well-tuned heuristic policies to ensure consistent data quality.
Figure 4: Trajectory smoothness comparison across simulated manipulation tasks. Lower values indicate smoother execution.
Baselines.
We compare our method against representative diffusion- and flow-based visuomotor policies. DP3 [33] serves as the primary point-cloud-based diffusion baseline, with iDP3 [32] and Reactive Diffusion Policy (RDP [28]) included as advanced variants, where the former employs a more expressive scene encoder, and the latter augments diffusion-based policies with an additional non-recurrent encoder. Due to environment constraints, RDP is evaluated without force or pressure sensing. We further include the original 2D RGB-based Diffusion Policy [5] as a canonical reference in visuomotor imitation learning, and a standard Flow Policy [35] that applies continuous flow matching directly to robotic action generation, serving as a representative flow-based control baseline.
Training and Implementation Details.
All methods condition on the two most recent observations and predict actions for the subsequent four steps. For each task, 30 expert demonstrations are collected, and performance is evaluated over 50 rollouts. RGB observations are cropped to $84\times 84$, and point clouds are downsampled using farthest point sampling to 512 points. Models are trained for 150 epochs using AdamW with a learning rate of $1\times 10^{-4}$, batch size 96, and EMA with decay rate 0.95. State and action inputs are normalized to [1,1][-1,1], with actions unnormalized before execution. All experiments are conducted on identical hardware with an Intel Core i7-14700KF CPU and an NVIDIA RTX 4090D GPU. In simulation, no additional execution-time sensory signals are provided to the decoder, and action decoding relies solely on the latent action representation.
Figure 5: Real-world experimental setup and observations. Left: Franka Emika Panda robot with a LEAP Hand and the visual sensing setup (global L515 and wrist-mounted D435). Right: Example multimodal observations (RGB images, point clouds, wrist views) and task objects used in real-world experiments.
IV-B2 Simulation Results Analysis
Quantitative real-world results are summarized in Table I and Fig. 4, evaluating trajectory smoothness, task success, and response time.
Trajectory Smoothness.
Fig. 4 reports trajectory smoothness across all simulated tasks. CoLA-Flow Policy achieves the lowest average smoothness score of $0.052$, reducing the metric by $51.4\%$, $69.4\%$, and $77.2\%$ relative to RDP, DP3, and Flow Policy, respectively. RDP exhibits lower smoothness than other diffusion-based baselines, indicating that latent action representations partially improve trajectory stability. However, its trajectories remain less smooth than those of CoLA-Flow Policy, as RDP does not adopt temporally continuous latent action modeling and does not explicitly structure its latent space to be continuous over time. By contrast, CoLA-Flow Policy performs flow matching over temporally coherent latent trajectories, consistently suppressing high-frequency oscillations.
Task Success Rate and Response Time.
Table I summarizes task success rates and inference latency across simulated benchmarks. CoLA-Flow Policy achieves the highest average success rate of $78.3\%$, outperforming Flow Policy ($17.3\%$) and DP3 ($12.6\%$), and improving over RDP ($73.0\%$) by $5.3\%$. While Flow Policy attains the lowest latency (6.1 ms), CoLA-Flow Policy incurs only a marginal increase (7.5 ms) and remains in the same real-time regime. Notably, CoLA-Flow Policy is approximately $7.5\times$ faster than DP3 (55.9 ms), demonstrating a favorable success–latency trade-off with minimal loss in responsiveness.
Overall, simulation results show that CoLA-Flow Policy achieves a favorable trade-off among smooth execution, fast inference, and high task success, which is essential for real-time robotic manipulation.
IV-C Real-World Experiments
IV-C1 Real-World Experimental Setup
Platform and Sensing.
Real-world experiments are conducted on a Franka Emika Panda arm with an interchangeable end-effector configuration, where the LEAP Hand can be replaced by a parallel gripper depending on the task (Fig. 5, left). A RealSense L515 global camera provides scene-level RGB-D observations, while a wrist-mounted RealSense D435 supplies local visual feedback. Both cameras operate at $640\times 480$ resolution and 30 fps. Depth images are converted to point clouds using calibrated intrinsics and extrinsics. In contrast to simulation, RGB images from the wrist-mounted camera are additionally used as execution-time conditioning inputs to the decoder, allowing low-level action realization to adapt to local visual variations during real-world execution.
Tasks.
We evaluate three real-world manipulation tasks: pick-and-place, performed with both the LEAP Hand and a parallel gripper to cover different manipulation modalities; peg-in-hole, which requires precise insertion; and obstacle-avoidance grasping, where a target object must be grasped while avoiding surrounding obstacles. Representative objects and environments are shown in Fig. 5 (top right).
Demonstrations and Baselines.
Expert demonstrations are collected via human teleoperation, with 30 demonstrations per task. We compare CoLA-Flow Policy against DP3 [33], Reactive Diffusion Policy (RDP) [28], and Flow Policy [35]. For RDP, tactile inputs are replaced with wrist camera images to ensure a fair comparison under visual-only settings.
Evaluation Protocol.
Each task is evaluated over 3 independent runs with different random seeds, each consisting of 10 trials (30 trials per task). Success rates are reported as mean $\pm$ standard deviation. Response time measures end-to-end policy inference latency, excluding sensor acquisition and environment stepping. All experiments are conducted on identical hardware with an Intel Core i9-14900KF CPU and an NVIDIA RTX 4080 GPU.
IV-C2 Real-World Results Analysis
Quantitative real-world results are summarized in Table II and Fig. 6, evaluating trajectory smoothness, task success, and response time.
TABLE II: Main results for real-world robot experiments. Success rates (%) are reported as mean $\pm$ standard deviation over 30 trials. Avg. Success denotes the average success rate across all tasks. Avg. Response Time measures end-to-end policy inference latency per control step.
Methods Pick & Place Pick & Place (Gripper) Peg-in-Hole Obstacle Avoidance Avg. Success Avg. Response Time
DP3 $66.7\pm 5.8$ $80.0\pm 5.8$ $35.0\pm 5.8$ $50.0\pm 10.0$ $57.9$ 29.29 $29\text{ ms}$
RDP $73.3\pm 5.8$ $93.3\pm 5.8$ $50.0\pm 0.0$ $63.3\pm 5.8$ $70.0$ $31.82\text{ ms}$
Flow Policy $60.0\pm 0.0$ $73.3\pm 5.8$ $30.0\pm 10.0$ $46.7\pm 5.8$ $52.5$ $\mathbf{6.54\text{ ms}}$
CoLA-Flow Policy $\mathbf{83.3\pm 5.8}$ $\mathbf{96.7\pm 5.8}$ $\mathbf{60.0\pm 0.0}$ $\mathbf{70.0\pm 5.8}$ $\mathbf{77.5}$ $8.59\text{ ms}$
Trajectory Smoothness.
Figure 6: Trajectory smoothness comparison across real-world manipulation tasks. Lower values indicate smoother execution.
As shown in Fig. 6, CoLA-Flow Policy achieves the lowest mean and maximum smoothness scores, indicating substantially reduced jerk and high-frequency oscillations in real-world execution. Compared to Flow Policy, the mean smoothness is reduced by over $93.7\%$, demonstrating robustness to sensing noise and actuation delays. Consistent with simulation results, RDP improves smoothness over other diffusion-based baselines, but remains less smooth than CoLA-Flow Policy due to the lack of explicit temporal smoothness constraints in the latent space.
This effect is further illustrated in Fig. 7, which shows joint trajectories over time under identical initial conditions. Flow Policy exhibits pronounced high-frequency oscillations, whereas CoLA-Flow Policy yields smoother, temporally coherent joint motions.
Figure 7: Comparison of real-world joint trajectories under identical initial conditions. Flow Policy exhibits high-frequency oscillations and abrupt variations, while CoLA-Flow Policy generates smoother and more temporally coherent trajectories, consistent with the quantitative smoothness metrics.
Task Success Rate and Response Time.
Table II reports task success rates and inference latency in real-world experiments. CoLA-Flow Policy achieves the highest average success rate of $77.5\%$, outperforming DP3 ($57.9\%$), RDP ($70.0\%$), and Flow Policy ($52.5\%$). While Flow Policy attains the lowest latency (6.54 ms), its unstable trajectories often lead to execution failures. CoLA-Flow Policy incurs only a modest increase in response time (8.59 ms) while achieving substantially higher task success. One contributing factor to the improved success of both RDP and CoLA-Flow Policy is the incorporation of wrist-mounted camera observations during action decoding, providing execution-time visual feedback for precise manipulation. CoLA-Flow Policy remains over $3\times$ faster than DP3 and RDP, indicating that significant real-world performance gains are achieved at minimal computational cost. Overall, real-world results confirm that CoLA-Flow Policy strikes a robust balance between inference efficiency, trajectory smoothness, and task success, supporting its suitability for stable real-world deployment.
IV-D Ablation Study
Figure 8: Ablation study on trajectory smoothness and task success rate in real-world manipulation tasks.
Fig. 8 summarizes ablation results in both simulation and real-world settings, analyzing the contributions of geometry-aware perception (GE-Flow) and latent action modelling (L-Flow). Across both domains, the two components exhibit complementary effects on task success and trajectory smoothness.
In simulation, geometry-aware encoding improves average success from $61\%$ to $64.3\%$ but slightly degrades trajectory smoothness, indicating that enhanced spatial perception alone does not resolve temporal instability. In contrast, latent action modelling substantially reduces the smoothness metric ($0.228\rightarrow 0.046$) while increasing success to $72.4\%$. Combining both components further improves success to $78.3\%$ while maintaining low oscillations.
A similar trend is observed in real-world experiments. While geometry-aware perception alone yields only marginal success gains ($52.5\%\rightarrow 54.2\%$), it noticeably increases trajectory oscillations, often leading to execution failures under real-world noise and actuation delays. Latent action modelling, by contrast, improves both metrics, increasing success to $71.7\%$ and reducing smoothness by over $95\%$. Combining both components achieves the best real-world performance, reaching a success rate of $77.5\%$ with minimal oscillations.
Overall, these results indicate that geometry-aware perception enhances spatial understanding, whereas stable manipulation critically relies on enforcing temporal coherence through latent action modelling, particularly in real-world settings.
V CONCLUSION
This work addresses real-time robotic manipulation under long-horizon and high-dimensional control, where existing generative policies struggle to jointly achieve expressive behavior modelling, fast inference, and stable execution, limiting their deployment on physical systems.
We propose CoLA-Flow Policy, a trajectory-level action generation framework that performs flow matching in a coherent latent action space. By mapping action sequences into a temporally coherent latent trajectory manifold and learning an explicit latent-space flow, CoLA-Flow Policy enables near single-step inference while substantially improving execution smoothness and stability. Operating on motion-level latent representations decouples global trajectory structure from low-level control noise, mitigating error amplification during generation and leading to more reliable long-horizon behaviours.
CoLA-Flow Policy further integrates geometry-aware point cloud conditioning and execution-time decoding modulation, allowing additional sensory signals to be incorporated without perturbing latent trajectory dynamics. This design enables richer multimodal execution while preserving fast response, as demonstrated by visual-conditioned decoding with a wrist-mounted camera in real-world experiments. Extensive simulation and real-robot results show smooth trajectories, high success rates, and rapid inference, outperforming both diffusion-based and raw action-space flow-based baselines.
Despite its effectiveness, the current system adopts relatively lightweight multimodal fusion and does not explicitly model complex contact dynamics. Future work will explore richer multimodal representations and contact-aware latent dynamics to extend CoLA-Flow Policy toward more challenging contact-rich manipulation.